determine the velocity of steam flowing in a circular pipe with a diameter of 4 inches and volume flow rate of 4 liters per second.


1
Expert's answer
2021-07-01T03:26:01-0400

"\frac{dQ}{dt}=Av\rightarrow (1)"

"A=\pi r^2"

r=2inch

"r=2\times254\times10^{-4}m"

"\frac{dQ}{dt}=4 liter/sec"

"\frac{dQ}{dt}=4\times10^{-3}m^3"


"A=\pi r^2=3.14\times(0.0508)^2=8.103\times10^{-3}m^2"

"v=\frac{\frac{dQ}{dt}}{A}=\frac{4\times10^{-3}}{8.103\times10^{-3}}=0.50m/sec"